# Training Neural Networks from Scratch with Parallel Low-Rank Adapters

### v1

@@ 1 Introduction
The escalating complexity of state-of-the-art deep learning models presents significant challenges not only in terms of computational demand but also in terms of memory and communication bandwidth. As these demands exceed the capacity of consumer-grade GPUs, training larger models requires innovative solutions. A prominent example for finetuning models is low-rank adaptation (Hu et al., 2022) that uses a low-rank parameterization of a deep neural network to reduce memory requirements for storing optimizer-state and gradient communication during training. The memory requirement was further reduced by quantizing model parameters (Dettmers et al., 2023). Such innovations have enabled the finetuning of large models, even on a single consumer-grade GPU.
@@
  • In Section 4 we combine LTE with federated averaging (McMahan et al., 2017), to demonstrate comparable performance to distributed pre-training, even with infrequent synchronization.
@@
Differences to existing works
Our work explores a new territory in the pre-training paradigm, leveraging both data and model parallelism. Our approach stores different copies of the LoRA parameters and is trained on different shards of the data distribution. This is in contrast with traditional federated learning, which replicates the same model across devices, and in-silo distributed training, which either divides the model across many devices for model parallelism or replicates the exact same model across devices for data parallelism. Our method enables distributed training with infrequent synchronizations while allowing for single-device inference. Several concurrent works have aimed at similar goals. We highlight a few notable works below:
ReLoRA (Lialin et al., 2023) sequentially trains and merges LoRA parameters back into the main weights. However, it does not yield comparable pre-training performance without initial full-parameter training. In contrast, our work uses parallel updates to match pre-training performance without bells and whistles.
FedLoRA (Yi et al., 2023) is designed to train LoRA parameters for finetuning within a federated learning framework. It involves training multiple LoRAs and averaging them into a single global LoRA. FedLoRA focuses on the distributed finetuning of LoRA parameters.
AdaMix (Wang et al., 2022) averages all MLPs in a Mixture of Experts (MOE) into a single MLP. AdaMix requires constant synchronization during the forward and backward passes. Our work requires no synchronization in the forward pass, and merging of the parameters can happen infrequently.
For an in-depth discussion of related works, see Section 5.
@@ 3 Method
To understand the conditions required to pre-train a model with LoRA, we first identify a specific scenario where standard training performance can be recovered using LoRA. This serves as a guide for developing our algorithm that retains the memory efficiency of LoRA.
@@ 3.1 Motivation: Multi-head merging perspective
Figure 3: LTE diagram: Our method is decomposed into 3 steps. (1) We parameterize the model with multiple LoRA heads and train them independently for $T$ iterations using different mini-batches sampled from the same (homogeneous) distribution. This results in overall update of $\delta_{\mathsf{lora}_{n}}(\mathbf{x})=-\eta\sum_{t}\nabla_{\mathsf{lora}_{n}}(\mathbf{x}[t])$ (2). Next, we accumulate the individual LoRA updates by averaging the heads $\Delta_{\mathsf{lora}}(\mathbf{x})=\frac{1}{N}\sum_{n}\delta_{\mathsf{lora}_{n}}(\mathbf{x})$. (3) The update is applied to the main weights, and the LoRA parameter $\mathbf{B}$ is reset. The optimization repeats with the new LoRA parameters.
@@ 3.2 LoRA soup: delayed LoRA merging
To further reduce the communication cost of LTE, we extend and combine the ideas of local updates (McMahan et al., 2017) and model-averaging (Wortsman et al., 2022; Yadav et al., 2023; Ilharco et al., 2023). Instead of merging every iteration, we allow the LoRA parameters to train independently of each other for a longer period before the merge operator. This is equivalent to using stale estimates of the LoRA parameters $\hat{\mathbf{W}}=\mathbf{W}+\frac{s}{N}\sum_{j\neq n}^{N}\mathbf{B}_{j}^{\prime}\mathbf{A}_{j}^{\prime}$ with ${}^{\prime}$ indicating a stale estimate of the parameters.
@@ 3.4 Implementation details
It is a common misconception that scaling $s$ has the same effect as tuning the learning rate $\eta$. During our experimentation, we were unable to yield comparable performance when using the standard value of $s$ (in the range of $1\sim 4$). Instead, we found using a large value of $s$ and a slightly lower learning rate $\eta$ to work the best. The standard practice is to set the scaling proportionately to the rank of the LoRA $s=\alpha/r$. This is done to automatically adjust for the rank (Hu et al., 2022). We use $\alpha=4096$ ($s=64$) and a learning rate of $\eta=2\cdot 10^{-4}$. It is worth noting that the learning rate does not scale linearly with $s$, and the scalar only affects the forward computation (Appendix B.1). The scalar $s$ modifies the contribution of the LoRA parameters in the forward pass which has a non-trivial implication on the effective gradient. Moreover, in Appendix B.2, we provide the effective update rule of LoRA updates and observe the emergence of the term $s$ that scales quadratically with the alignment of $\mathbf{B}$ and $\mathbf{A}$. Since $s$ is large relative to the learning rate, it may have a non-negligible effect on the dynamics.
Input: Dataset $\mathcal{D}_{\mathsf{train}}$, model $\mathcal{F}$, loss function $\mathcal{L}$     parameters $\Theta=\{\mathbf{W}_{0},\dots,\mathbf{W}_{L}\}$,     merge scalar $s$, num workers $N$, merge iter $T$   while not converged do        (Optional) quantize $\Theta$. Keep high-precision copy        (in parallel) for each worker $n$ do              if LoRA not initialized then                    $\mathcal{B}_{n},\mathcal{A}_{n}\leftarrow\mathbf{lora\_parameterize}(\mathcal{F})$             else                    (Optional) reset parameters $\mathcal{B}_{n}$ to zero             Optimize $\mathcal{B}_{n},\mathcal{A}_{n}$ for $T$ iterations by minimizing $\mathbb{E}_{\mathbf{x},\mathbf{y}\sim\mathcal{D}_{\mathsf{train}}}\left[\mathcal{L}(\mathcal{F}(\mathbf{x}),\mathbf{y})\right]$        # Synchronize by communicating LoRA parameters. for each worker $n$ do              for $\mathbf{B}_{n},\mathbf{A}_{n}$ in $\mathcal{B}_{n},\mathcal{A}_{n}$ do                    Merge LoRA params $\mathbf{W}_{n}\leftarrow\mathbf{W}_{n}+\frac{s}{n}\mathbf{B}_{n}\mathbf{A}_{n}$ Algorithm 1 LoRA-the-Explorer (LTE)
@@ 4 Experiments
Disclaimer: During the preparation of the code release, we found that the transformer experiments used a scaling factor of $1/\sqrt{d_{out}}$ instead of the standard scaling $1/\sqrt{d_{out}/n_{attn}}$ (Vaswani et al., 2017). Hence using LTE with the standard scaling would require a different set of hyper-parameters than the ones reported in Appendix A.1. We will revise the hyper-parameters in the next revision.
@@ 4.1 Iterative LoRA Merging
We find that the merge iteration of $T=10$ is still stable when using batch size $4096$. With higher $T$ values, additional training may be required to achieve comparable performance (Stich, 2019; Wang & Joshi, 2021; Yu et al., 2019). Our initial efforts to improve merging using methods such as (Yadav et al., 2023) did yield better results. Nonetheless, we believe with increased merge iteration, smarter merging techniques may be necessary. Existing literature in federated learning and linear-mode connectivity/model-averaging may provide insights in designing better merging criteria.
@@ 4.2 LoRA parameter alignment
LTE Heads Rank Merge Test loss $\downarrow$ Test acc $\uparrow$ - - - - 1.78 71.97 $\mathsf{head}$ 2 8 10 2.31 54.68 8 8 10 2.35 54.88 32 8 10 2.26 58.21 2 64 10 1.86 69.82 8 64 10 1.84 71.97 32 64 10 1.79 73.73 $\mathsf{rank}$ 32 8 10 2.66 44.00 32 16 10 2.12 60.35 32 32 10 1.81 71.20 32 64 10 1.79 73.73 32 128 10 1.78 73.54 $\mathsf{merge}$ 32 64 5 1.87 70.67 32 64 10 1.79 73.73 32 64 20 1.97 66.80 32 64 50 1.99 65.43 32 64 100 2.10 61.04 Table 1: LTE ablation for ViT-S trained ImageNet100: For fixed cumulative training epoch of $1200$, we vary the number of heads, rank, and merge iteration of our method. More heads require longer cumulative training samples to converge; see Figure 8.
@@ 4.3 Ablation study: the effect of LoRA heads, rank, and merge iteration
We systematically evaluate the effects of varying the number of LoRA heads, rank, and merge iteration on model performance for ImageNet100 in Table 1. Our findings indicate a monotonic improvement in performance with an increased number of heads and ranks. Conversely, extending the merge iteration negatively impacts performance. As in the case of least-squares regression, we found excessive merging to hurt model accuracy. With a large enough rank and head, we found the model to converge to better test accuracy, even if the test loss was similar. We hypothesize the averaging of the LoRA heads has a regularization effect similar to that model ensembling.
We use ViT-S as the primary architecture for analysis, which has a hidden dimension of $384$ and an MLP dimension of $1536$. We find that setting the product of the number of heads and the rank of the LoRA larger than the largest dimension of the model serves as a good proxy for configuring LTE. For example, using $32$ heads with $r=64$ results in $2048>1536$. However, when it comes to increasing the number of heads rather than rank, we noticed longer training iterations were required to achieve comparable performance. We discuss a potential cause of the slowdown in convergence in the preceding section.
@@ 5 Related works
The use of LoRA has garnered considerable attention in recent research. While our work focuses on using adapters to pre-train model from scratch, majority of the works have focused on enhancing finetuning processes through these adapters (Chavan et al., 2023; Zhang et al., 2023b), while others aim to diminish computational requirements (Zhang et al., 2023a) or to offset part of the pre-training computation (Lialin et al., 2023). Moreover, motivated by (Wortsman et al., 2022)(Wang et al., 2022) proposes the use mixture-of-expert for parameter efficient finetuning and uses averaging for efficient inference. Various forms of adapters have been proposed in prior research, each serving specific applications. Additive adapters (Zhang et al., 2021) augment the model size for inference, and hence the community has turned towards linear adapters either in the form residual connections (Cai et al., 2020) or affine parameters in batch-norm, (Bettelli et al., 2006; Mudrakarta et al., 2018). Adapters have been applied for natural language processing (Houlsby et al., 2019; Stickland & Murray, 2019), video (Yang et al., 2024b; Xing et al., 2023), computer vision (Sax et al., 2020; Zhang & Agrawala, 2023; Chen et al., 2023), incremental learning (Rosenfeld & Tsotsos, 2018), domain adaptation (Rebuffi et al., 2018), and vision-language tasks (Gao et al., 2023; Radford et al., 2021; Sung et al., 2022), text-to-vision generative models (Mou et al., 2023), and even perceptual learning (Fu et al., 2023).
@@
Our work has relevance to both distributed and federated learning paradigms, wherein each head is conceptualized as a distinct computational device. Federated learning addresses various topics, including low-compute devices, high-latency training, privacy, and both cross and in-silo learning, as comprehensively discussed in (McMahan et al., 2017; Wang et al., 2021). Communication efficiency serves as a cornerstone in both distributed and federated learning. Techniques such as local steps have been employed to mitigate communication load (McMahan et al., 2017; Lin et al., 2020; Povey et al., 2014; Smith et al., 2018; Su & Chen, 2015; Zhang et al., 2016). These methods defer the averaging of weights to specific optimization steps, thus alleviating the communication cost per iteration. The effectiveness of decentralized training has been studied in (Lian et al., 2017; Koloskova et al., 2019, 2020; Coquelin et al., 2022). Traditionally, activation computations have dominated the computational load. However, the advent of gradient checkpointing (Chen et al., 2016) and reversible gradient computation (Gomez et al., 2017; Mangalam et al., 2022) has shifted the training process toward being increasingly parameter-bound. Techniques such as gradient or weight compression also seek to reduce the communication burden (Lin et al., 2018; Aji & Heafield, 2017; Wen et al., 2017). Combining models in federated learning is often credited to FedAvg (McMahan et al., 2017). Numerous studies explore the use of weighted averaging to improve convergence speed (Li et al., 2020b). Since then, many works have tried to use probabilistic frameworks to understand and improve merging (Hsu et al., 2019; Wang et al., 2020; Reddi et al., 2021). The conditions for optimal merging is still an open question, with recent efforts to improve updating with stale parameters has been explored in (Chen et al., 2022). Server momentum and adaptive methods constitute another active area of research where macro synchronization steps may be interpreted as gradients, thus facilitating bi-level optimization schemes (Hsu et al., 2019; Wang et al., 2020; Reddi et al., 2021). Initial efforts to employ federated learning with large models have been made. (Yuan et al., 2022) examined the cost models for pre-training Language Learning Models (LLMs) in a decentralized configuration. (Wang et al., 2023) suggested the utilization of compressed sparse optimization methods for efficient communication.
@@
Linear mode connectivity (Garipov et al., 2018) pertains to the study of model connectivity. Deep models are generally linearly disconnected but can be connected through nonlinear means (Freeman & Bruna, 2017; Draxler et al., 2018; Fort & Jastrzebski, 2019). Under the same initialization, linear paths with constant energy exist in trained models (Nagarajan & Kolter, 2019; Frankle et al., 2020; Wortsman et al., 2022). For models with different initializations, parameter permutations can be solved to align them linearly (Brea et al., 2019; Tatro et al., 2020; Entezari et al., 2021; Simsek et al., 2021). Following this line of research, numerous works have delved into model averaging and stitching. Where averaging of large models has shown to improve performance (Wortsman et al., 2022; Ainsworth et al., 2023; Stoica et al., 2024; Jordan et al., 2023). Model stitching (Lenc & Vedaldi, 2015) has also shown to yield surprising transfer capabilities (Moschella et al., 2023). This idea is conceptually related to optimal averaging in convex problems (Scaman et al., 2019) and the “Anna Karenina” principle where successful models converge to similar solutions (Bansal et al., 2021). The effectiveness of averaging models within ensembles is well-established (Huang et al., 2017; Izmailov et al., 2018; Polyak & Juditsky, 1992). Utilizing an average model as a target has also been investigated (Tarvainen & Valpola, 2017; Cai et al., 2021; Grill et al., 2020; Jolicoeur-Martineau et al., 2023; Wortsman et al., 2023).

### v2

@@ 1 Introduction
The escalating complexity of state-of-the-art deep learning models presents significant challenges not only in terms of computational demand but also in terms of memory and communication bandwidth. As these demands exceed the capacity of consumer-grade GPUs, training larger models requires innovative solutions. A prominent example of finetuning models is low-rank adaptation (Hu et al., 2022) that uses a low-rank parameterization of a deep neural network to reduce memory requirements for storing optimizer-state and gradient communication during training. The memory requirement was further reduced by quantizing model parameters (Dettmers et al., 2023). Such innovations have enabled the finetuning of large models, even on a single consumer-grade GPU.
@@
  • In Section 4, we combine LTE with federated averaging (McMahan et al., 2017) to demonstrate comparable performance to distributed pre-training, even with infrequent synchronization.
@@
  • We provide discussion of related works in Section 5.
@@ 3 Method
To understand the conditions required to pre-train a model with LoRA, we first identify a specific scenario where standard training performance can be recovered using LoRA. This serves as a guide for developing our algorithm that retains LoRA’s memory efficiency.
@@ 3.1 Motivation: Multi-head merging perspective
Figure 3: LTE diagram: Our method is decomposed into 3 steps. (1) We parameterize the model with multiple LoRA heads and train them independently for $T$ iterations using different mini-batches sampled from the same (homogeneous) distribution. This results in overall update of $\delta_{\mathsf{lora}_{n}}(\mathbf{x})=-\eta\sum_{t}\nabla_{\mathsf{lora}_{n}}(\mathbf{x}[t])$ (2). Next, we accumulate the individual LoRA updates by averaging the heads $\Delta_{\mathsf{lora}}(\mathbf{x})=\frac{1}{N}\sum_{n}\delta_{\mathsf{lora}_{n}}(\mathbf{x})$. (3) The update is applied to the main weights, and the LoRA parameter $\mathbf{B}$ is reset. The optimization repeats with the new LoRA parameters. LTE resembles the distributed model development paradigm first proposed by Kandpal et al. (2023).
@@ 3.2 LoRA soup: delayed LoRA merging
To further reduce the communication cost of LTE, we extend and combine the ideas of local updates (McMahan et al., 2017) and model-averaging (Wortsman et al., 2022; Yadav et al., 2023; Ilharco et al., 2023). Instead of merging every iteration, we allow the LoRA parameters to train independently of each other for a longer period before the merge operator. This is equivalent to using stale estimates of the LoRA parameters $\hat{\mathbf{W}}=\mathbf{W}+\frac{s}{N}\sum_{j\neq n}^{N}\mathbf{B}_{j}^{\prime}\mathbf{A}_{j}^{\prime}$ with indicating a stale estimate of the parameters.
@@ 3.4 Implementation details
It is a common misconception that scaling $s$ has the same effect as tuning the learning rate $\eta$. During our experimentation, we were unable to yield comparable performance when using the standard value of $s$ (in the range of $1\sim 4$). Instead, we found using a large value of $s$ and a slightly lower learning rate $\eta$ to work the best. The standard practice is to set the scaling proportionately to the rank of the LoRA $s=\alpha/r$. This is done to automatically adjust for the rank (Hu et al., 2022). We use $\alpha=4096$ ($s=64$) and a learning rate of $\eta=2\cdot 10^{-4}$. It is worth noting that the learning rate does not scale linearly with $s$, and the scalar only affects the forward computation (Appendix B.1). The scalar $s$ modifies the contribution of the LoRA parameters in the forward pass, which has a non-trivial implication on the effective gradient. Moreover, in Appendix B.2, we provide the effective update rule of LoRA updates and observe the emergence of the term $s$ that scales quadratically with the alignment of $\mathbf{B}$ and $\mathbf{A}$. Since $s$ is large relative to the learning rate, it may have a non-negligible effect on the dynamics.
Input: Dataset $\mathcal{D}_{\mathsf{train}}$, model $\mathcal{F}$, loss function $\mathcal{L}$
    parameters $\Theta=\{\mathbf{W}_{0},\dots,\mathbf{W}_{L}\}$,
    merge scalar $s$, num workers $N$, merge iter $T$
while not converged do
       (Optional) quantize $\Theta$. Keep high-precision copy
       (in parallel) for each worker $n$ do
             if LoRA not initialized then
                   $\mathcal{B}_{n},\mathcal{A}_{n}\leftarrow\mathbf{lora\_parameterize}(\mathcal{F})$
            else
                   (Optional) reset parameters $\mathcal{B}_{n}$ to zero
            Optimize $\mathcal{B}_{n},\mathcal{A}_{n}$ for $T$ iterations by minimizing $\mathbb{E}_{\mathbf{x},\mathbf{y}\sim\mathcal{D}_{\mathsf{train}}}\left[\mathcal{L}(\mathcal{F}(\mathbf{x}),\mathbf{y})\right]$
       # Synchronize by communicating LoRA parameters.
       for each worker $n$ do
             for $\mathbf{B}_{n},\mathbf{A}_{n}$ in $\mathcal{B}_{n},\mathcal{A}_{n}$ do
                   Merge LoRA params $\mathbf{W}_{n}\leftarrow\mathbf{W}_{n}+\frac{s}{n}\mathbf{B}_{n}\mathbf{A}_{n}$
Algorithm 1 LoRA-the-Explorer (LTE)
@@ 4 Experiments
Disclaimer: During the preparation of the code release, we found that the transformer experiments used a scaling factor of $1/\sqrt{d_{out}}$ instead of the standard scaling $1/\sqrt{d_{out}/n_{attn}}$ (Vaswani et al., 2017). Hence, using LTE with the standard scaling would require a different set of hyper-parameters than the ones reported in Appendix A.1. We will revise the hyper-parameters in the next revision.
@@ 4.1 Iterative LoRA Merging
We find that the merge iteration of $T=10$ is still stable when using batch size $4096$. With higher $T$ values, additional training may be required to achieve comparable performance (Stich, 2019; Wang & Joshi, 2021; Yu et al., 2019). Our initial efforts to improve merging using methods such as (Yadav et al., 2023) did yield better results. Nonetheless, we believe with increased merge iteration, smarter merging techniques may be necessary. Existing literature in federated learning and linear-mode connectivity/model-averaging may provide insights into designing better merging criteria.
@@ 4.2 LoRA parameter alignment
LTE Heads Rank Merge Test loss $\downarrow$ Test acc $\uparrow$
- - - - 1.78 71.97
$\mathsf{head}$ 2 8 10 2.31 54.68
8 8 10 2.35 54.88
32 8 10 2.26 58.21
2 64 10 1.86 69.82
8 64 10 1.84 71.97
32 64 10 1.79 73.73
$\mathsf{rank}$ 32 8 10 2.66 44.00
32 16 10 2.12 60.35
32 32 10 1.81 71.20
32 64 10 1.79 73.73
32 128 10 1.78 73.54
$\mathsf{merge}$ 32 64 5 1.87 70.67
32 64 10 1.79 73.73
32 64 20 1.97 66.80
32 64 50 1.99 65.43
32 64 100 2.10 61.04
Table 1: LTE ablation for ViT-S trained ImageNet100: For fixed cumulative training epoch of $1200$, we vary the number of heads, rank, and merge iteration of our method. More heads require longer cumulative training samples to converge; see Figure 9.
@@ 4.3 Ablation study: the effect of LoRA heads, rank, and merge iteration
We systematically evaluate the effects of varying the number of LoRA heads, rank, and merge iteration on model performance for ImageNet100 in Table 1. Our findings indicate a monotonic improvement in performance with an increased number of heads and ranks. Conversely, extending the merge iteration negatively impacts performance. As in the case of least-squares regression, we found excessive merging to hurt model accuracy. With a large enough rank and head, we found the model to converge to better test accuracy, even if the test loss was similar. We hypothesize that the averaging of the LoRA heads has a regularization effect similar to those seen in model ensembling.
We use ViT-S as the primary architecture for analysis, which has a hidden dimension of $384$ and an MLP dimension of $1536$. We find that setting the product of the number of heads and the rank of the LoRA larger than the largest dimension of the model serves as a good proxy for configuring LTE. For example, using $32$ heads with $r=64$ results in $2048>1536$. However, when it comes to increasing the number of heads rather than rank, we noticed longer training iterations were required to achieve comparable performance. In the preceding section, we discuss a potential cause of the slowdown in convergence.
@@ 5 Related works
The use of LoRA has received considerable attention in recent literature. While we focused on using LoRA to pre-train models from scratch, prior works have mainly focused on finetuning (Chavan et al., 2023; Zhang et al., 2023b). The flexibility of LoRA has had a broad reach in different aspects of model training, from reducing computational requirements (Zhang et al., 2023a) to offsetting part of the pre-training computation (Lialin et al., 2023).
LoRA is just one type of adapter, and various forms of adapters have been explored. Parallel additive adapters (Zhang et al., 2021) augment existing models with additional parameters. Linear adapters have also been common, either in the form of residual connections (Cai et al., 2020) or affine parameters in normalization layers (Bettelli et al., 2006; Mudrakarta et al., 2018).
Adapters have been applied to numerous tasks: natural language processing (Houlsby et al., 2019; Stickland & Murray, 2019), video (Yang et al., 2024b; Xing et al., 2023), computer vision (Sax et al., 2020; Zhang & Agrawala, 2023; Chen et al., 2023), incremental learning (Rosenfeld & Tsotsos, 2018), domain adaptation (Rebuffi et al., 2018), and vision-language tasks (Gao et al., 2023; Radford et al., 2021; Sung et al., 2022), text-to-vision generative models (Mou et al., 2023), and even perceptual learning (Fu et al., 2023).
@@
Our work shares relevance to both distributed and federated learning paradigms, where the LoRA head in our work can be conceptualized as a distinct computational device. Our work shares many motivations for federated learning, including low-compute devices, high-latency training, privacy, and cross and in-silo learning. See (McMahan et al., 2017; Wang et al., 2021) for a comprehensive discussion.
Communication efficiency serves as a cornerstone in both distributed and federated learning. Techniques such as local steps have been employed to mitigate communication load (McMahan et al., 2017; Lin et al., 2020; Povey et al., 2014; Smith et al., 2018; Su & Chen, 2015; Zhang et al., 2016). These methods defer the averaging of weights to specific optimization steps, thus reducing the communication cost per iteration. The effectiveness of decentralized training has been studied in (Lian et al., 2017; Koloskova et al., 2019, 2020; Coquelin et al., 2022).
Traditionally, activation computations have dominated the computational load. However, the advent of gradient checkpointing (Chen et al., 2016) and reversible gradient computation (Gomez et al., 2017; Mangalam et al., 2022) has shifted the training process toward being parameter-bound. Techniques such as gradient or weight compression also seek to reduce the communication burden (Lin et al., 2018; Aji & Heafield, 2017; Wen et al., 2017).
Combining models in federated learning is often credited to FedAvg (McMahan et al., 2017). Numerous studies explore the use of weighted averaging to improve convergence speed (Li et al., 2020b). Since then, many works have tried to use probabilistic frameworks to understand and improve merging (Hsu et al., 2019; Wang et al., 2020; Reddi et al., 2021). The conditions for optimal merging is still an open question, with recent efforts to improve updating with stale parameters (Chen et al., 2022).
Server momentum and adaptive methods constitute another active area of research where macro synchronization steps are interpreted as “gradients”, allowing one to prescribe a bi-level optimization scheme (Hsu et al., 2019; Wang et al., 2020; Reddi et al., 2021).
Initial efforts have been made to use federated learning with large models. (Yuan et al., 2022) examined the cost models for pre-training Language Learning Models (LLMs) in a decentralized configuration. (Wang et al., 2023) suggested the utilization of compressed sparse optimization methods for efficient communication. Our work is also closely related to (Douillard et al., 2023), a concurrent work that explores the idea of distributing computing across smaller server farms while maintaining low communication costs.
@@
Linear mode connectivity (Garipov et al., 2018) studies the phenomena of why and when models are smoothly connected (Freeman & Bruna, 2017; Draxler et al., 2018; Fort & Jastrzebski, 2019). Under the same initialization, models have been shown to have a linear path with constant energy (Nagarajan & Kolter, 2019; Frankle et al., 2020; Wortsman et al., 2022). For models with different initializations, parameter permutations can be solved to align them linearly (Brea et al., 2019; Tatro et al., 2020; Entezari et al., 2021; Simsek et al., 2021).
Following this line of research, numerous works have also explored model averaging and stitching. Where averaging of large models has shown to improve performance (Wortsman et al., 2022; Ainsworth et al., 2023; Stoica et al., 2024; Jordan et al., 2023; Wortsman et al., 2023). Model stitching (Lenc & Vedaldi, 2015) has also shown to yield surprising transfer capabilities (Moschella et al., 2023). This idea is conceptually related to optimal averaging in convex problems (Scaman et al., 2019) and the “Anna Karenina” principle where successful models converge to similar solutions (Bansal et al., 2021). This phenomenon could help explain the success of averaging multiple LoRAs into a single global LoRA (Yi et al., 2023) and averaging MoE MLP layers into a single weight at inference (Wang et al., 2022).